OpenStudy (anonymous):

How do you solve y^4-5y^2+4=0. Please show work

OpenStudy (anonymous):

y^4-5y^2+4=0 (y^2-4)(y^2-1)=0 (y+2)(y-2)(y+1)(y-1)=0 y=-2, 2, -1, 1

OpenStudy (anonymous):

thank-you

OpenStudy (anonymous):

WELCOME! XD